son is 4 times younger than dad. son difference to dad is 27 how old is the son and dad?
Answer by nihar*2013(34)

Suppose age of father is x years and age of son is y years
son is 4 times younger than dad
x = 4y
x-4y=0---------(1)
x-y=27----------(2)
- + -
____________
-3y= -27
y=9
x-y=27
x-9=27
x=27+9
x=36 years.

Dad is 36 years old son is 9
Thanks